OpenStudy (nuccioreggie):

A set of equations is given below: Equation A: c = 2d + 1 Equation B: c = 3d + 5 Which of the following is a step that can be used to find the solution to the set of equations? 2d + 1 = 3d + 5 2d = 3d + 5 2d + 1 = 3d 2d + 5 = 3d + 1

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (mayankdevnani):

what do you think ?

OpenStudy (nuccioreggie):

c

OpenStudy (mayankdevnani):

how ?

OpenStudy (mayankdevnani):

you know that equation A = equation B right ?

OpenStudy (nuccioreggie):

6 = 4 + 2 6 = 3 + 3 6 = 6 always so 4 + 2 = 3 + 3

OpenStudy (mayankdevnani):

you are correct ! but its not the way to proceed the question !

OpenStudy (mayankdevnani):

do you agree that `c=c` ?

OpenStudy (nuccioreggie):

no so the answer would be D

OpenStudy (nuccioreggie):

i ment B sorry

OpenStudy (mayankdevnani):

nope ! try it again man someething is wrong in option D

OpenStudy (nuccioreggie):

wooaah lol im stupid

OpenStudy (mayankdevnani):

lol..

OpenStudy (mayankdevnani):

option A is correct !

OpenStudy (mayankdevnani):

because equation A = equation B

OpenStudy (mayankdevnani):

so option A proves that above statement